OpenStudy (anonymous):

Find the centroid (x¯,y¯) of the region bounded by: y=3x^2+3x, y=0, x=0, and x=4

OpenStudy (irishboy123):

https://i.gyazo.com/34d3aa81faa7d27ef7750b8308bb53a7.png \[\bar x = \dfrac{\int_{0}^{4} \; dx \quad xy}{\int_{0}^{4} \; dx \quad y }\]

OpenStudy (anonymous):

or \[x =\frac{ 1 }{ A } \int\limits_{0}^{4} xf(x) dx\] and \[y = \frac{ 1 }{ A } \int\limits_{0}^{4} \frac{ 1 }{ 2 } (f(x))^{2}\] where \[A = \int\limits_{0}^{4} f(x) dx\]